### v3

# Semi-supervised Underwater Image Enhancement Using A Physics-Aware Triple-Stream Network

Abstract
Underwater images normally suffer from degradation due to the transmission medium of water bodies. Both traditional prior-based approaches and deep learning-based methods have been used to address this problem. However, the inflexible assumption of the former often impairs their effectiveness in handling diverse underwater scenes, while the generalization of the latter to unseen images is usually weakened by insufficient data. In this study, we leverage both the physics-based Image Formation Model (IFM) and deep learning techniques for Underwater Image Enhancement (UIE). To this end, we propose a novel Physics-Aware Triple-Stream Underwater Image Enhancement Network, i.e., PATS-UIENet, which comprises a Direct Signal Transmission Estimation Steam (D-Stream), a Backscatter Signal Transmission Estimation Steam (B-Stream) and an Ambient Light Estimation Stream (A-Stream). This network fulfills the UIE task by explicitly estimating the degradation parameters of a revised IFM. We also adopt an IFM-inspired semi-supervised learning framework, which exploits both the labeled and unlabeled images, to address the issue of insufficient data. To our knowledge, such a physics-aware deep network and the IFM-inspired semi-supervised learning framework have not been used for the UIE task before. Our method performs better than, or at least comparably to, sixteen baselines across six testing sets in the degradation estimation and UIE tasks. These promising results should be due to the fact that the proposed method can not only model the degradation but also learn the characteristics of diverse underwater scenes. [^1] The models and code will be published on the acceptance of the paper.
Index Terms:
Underwater Image Enhancement (UIE), Underwater Image Processing, Image Formation Model (IFM), Deep learning, Semi-supervised Learning.
I Introduction
The images captured in the underwater environment play important roles in ocean exploration. However, these images normally suffer from different degradation, due to the wavelength-dependent light absorption and light scattering caused by the underwater transmission medium. In [1], the Image Formation Model (IFM) can be expressed as:
$I^{c}(x)=J^{c}(x)t^{c}(x)+(1-t^{c}(x))A^{c},$ (1)
where $I^{c}(x)$ ($c\in\{R,G,B\}$) is a degraded image, $J^{c}(x)$ is the underlying clean image, $t^{c}(x)=exp(-\beta^{c}d(x))$ denotes the transmission maps, $\beta^{c}$ is the attenuation factor for each channel, $d(x)$ is the scene distance map and $A^{c}$ is the ambient light. (See Fig. 1 for examples). In this context, $J^{c}(x)t^{c}(x)$ describes the distance-dependent color distortion, while $(1-t^{c}(x))A^{c}$ represents the scattering of the ambient light which decreases the contrast and lessens the visibility of the image. Despite the IFM had been widely used, its flaws in accurately modeling the degradation process was highlighted [2].
Figure 1: Examples of a degraded underwater image, the corresponding ambient light image $A^{c}$ and three transmission maps $t^{c}(x)$ ($c\in\{R,G,B\}$) estimated, and the enhanced image produced using the estimated parameters of the IFM [1].
Considering the degradation may interfere with downstream tasks, Underwater Image Enhancement (UIE) is key to underwater machine vision. Due to the complicated degradation process, this task is challenging. Many traditional methods [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] have been proposed, to address the challenge. These methods can be treated as the prior-based methods because they typically rely on a certain form of prior information. Some prior-based methods [3, 4, 5] aimed to estimate the parameters of the IFM. However, they normally struggle with adapting to diverse underwater scenes or estimating valid parameters when the scene violates the prior assumption [5]. On the other hand, the other prior-based methods [11, 12, 13, 14, 15] directly enhance the quality of underwater images without explicitly modeling the degradation process. Due to the lack of the knowledge of underwater scenes, they cannot effectively enhance the images captured in the complex scenes and often introduce artifacts.
In contrast, deep learning-based UIE approaches [18, 19, 20, 21, 22, 23, 24, 25, 26] aim to directly learn from different types of degraded images. Therefore, the problem with the limited static priors can be alleviated. Ideally, a large number of real-world labeled training images should be utilized. However, it is difficult to collect such a large labeled underwater data set in practice. Alternatively, the best enhanced image can be selected as the reference (ground-truth) by observers from the images processed by a set of UIE algorithms [22, 27]. Nevertheless, this sort of data sets cannot represent diverse underwater scenes due to the limited size. In this case, the deep UIE model trained may overfit the training data.
The synthetic images generated using the IFM [28, 18] or a Generative Adversarial Network (GAN) [29, 30, 31, 20, 21] were employed in order to address the problem with the lack of labeled data. However, the utilization of these images to simulate real-world underwater degradation processes is struggling. Transfer learning techniques [30, 19] were also applied for that purpose, which usually used generative models to map degraded images to the clearer images. But they are difficult to train and may produce unstable results. Recently, promising results were derived using zero-shot learning approaches [32, 33] by directly learning degradation parameters from unlabeled degraded images.
To effectively model the degradation process, we propose a Physics-Aware Triple-Stream Underwater Image Enhancement Network, i.e., PATS-UIENet, by injecting the physical principle into the deep neural network. This network performs the UIE task by explicitly estimating the degradation parameters of a revised IFM [2] using a Direct Signal Transmission Estimation Steam (D-Stream), a Backscatter Signal Transmission Estimation Steam (B-Stream) and an Ambient Estimation Stream (A-Stream). These parameters are fed into the revised IFM and the result is the image that the PATS-UIENet enhances.
We further adopt an IFM-inspired semi-supervised learning framework, including a bi-directional supervised scheme and an unsupervised scheme, to overcome the limitation of insufficient data. Compared with the unsupervised methods [32, 33], the framework can be more easily trained due to the facilitation from the supervised scheme. Since the unsupervised scheme uses additional unlabeled images while the bi-directional supervised scheme exploits labeled images more effectively, it can be better trained than the supervised approaches [18, 19, 22, 23, 24, 25, 26, 34]. As a result, the generalization of the model trained using our framework is stronger than that trained using a supervised or unsupervised method.
To our knowledge, this study makes the first effort to jointly apply the physics-aware deep network and the IFM-inspired semi-supervised learning technique to the UIE task. Our main contributions can be summarized as threefold.
  • We propose a physics-aware triple-stream network for the UIE task by explicitly estimating the degradation parameters of the revised IFM. Such an IFM-motivated network has not been explored for the UIE task before.
  • An IFM-inspired semi-supervised learning framework is adopted, to overcome the challenge of insufficient data. This framework exploits the advantages of both the supervised and unsupervised methods.
  • We conduct a series of comparative experiments on five real-world underwater testing sets along with sixteen baselines. These results provide the community with a set of benchmarks.
The remainder of this paper is organized as follows. The related literature is reviewed in Section II. We introduce the proposed method in Section III. The experimental settings and results are reported in Sections IV and V respectively. Finally, our conclusion is drawn in Section VI.
II Related Work
II-A Prior-Based Methods
To recover clear underwater images, many methods [3, 4, 5, 6, 7, 8, 10] used the prior information to estimate the degradation parameters of the IFM. Although the Dark Channel Prior (DCP) [3] was initially proposed for image dehazing, some researchers [6, 7] employed it for underwater image enhancement, due to the similarity between the degradation processes resulted from the foggy weather and underwater environment. However, the original DCP usually yielded erroneous estimations. Therefore, more studies were performed to improve it, such as Underwater Dark Channel Prior (UDCP) [4] and Generalization of the Dark Channel Prior (GDCP) [5]. Moreover, the Haze Line Prior (HL) [8] was utilized in some studies [10, 35]. Despite these methods were designed on top of the IFM, they normally made a rigid assumption about the underwater environment, which restricted the application of them to a specific underwater scenario.
On the other hand, some approaches used the more general priors [36, 11, 12, 13, 16, 14, 15] to enhance the quality of underwater images without taking the IFM into account, including Histogram Equalization (HE) [36, 11], Retinex-based methods [13, 16] and image fusion techniques [17, 37]. These approaches usually enhanced the contrast and produced more color-balanced results. However, they probably struggle with processing globally inhomogeneous degraded images and may introduce artifacts, such as halos and color casts, due to the lack of the knowledge of underwater scenes.
II-B Learning-Based Methods
Thanks to the powerful representation learning ability and large-scale training data [38], deep learning-based methods greatly boosted the development of computer vision in both high-level [39, 40, 41] and low-level vision tasks [42, 43, 44]. However, the application of deep learning to underwater image enhancement is much less than other tasks. This dilemma should be attributed to the difficulty on collecting a large number of underwater images and the corresponding clean counterparts. Some deep learning methods [28, 45] used the fake underwater images, synthesized based on terrestrial images, to train a network for degradation parameter estimation. Since the synthesis algorithms were relatively simple, they usually could not simulate diverse underwater scenes. Besides, the content of the terrestrial images was significantly different from underwater scenes. As a result, those methods usually encountered the domain-shift problem when they were applied to real underwater images.
To overcome this problem, some labeled real-world underwater image data sets, e.g., UIEB [22] and SUIM-E [27], were collected. Regarding a degraded image, a set of enhancement methods were applied. The most visually pleasant result was manually picked out as the reference of the degraded image. A UIE model can be trained using these data sets without considering the IFM. For example, the WaterNet [22] used CNNs to learn fusion weights for the purpose of fusing the enhanced results produced by the White Balance [17], HE and Gamma Correction techniques. To perform the UIE task, the UColor [23] method combined the RGB, HSV and Lab color spaces with the guidance of the IFM parameters estimated using the GDCP [5]. Since the training data sets were relatively small, these methods usually overfitted the training data and poorly generalized to unseen images. In addition, they could not explicitly utilize the information provided by the IFM.
In contrast to the lack of labeled underwater images, it is easier to collect unlabeled underwater images. Existing studies [19, 46] sourced real-world underwater images and categorized them according to the degree of degradation. Unsupervised learning approaches were developed on top of these data by learning the mapping from severely degraded images to slightly degraded images. However, these approaches were hard to train and tended to produce unstable results. Recently, a semi-supervised UIE method [47] was developed. Nevertheless, this method did not consider physical principles.
The above-mentioned studies either are not robust to diverse underwater scenes, or lack the knowledge of underwater images. We are hence motivated to exploit both the IFM-inspired semi-supervised learning technique and the physics-aware deep neural network, to address these issues.
III Methodology
Considering the importance of physical principles to the UIE task, we propose a Physics-Aware Triple-Stream Underwater Image Enhancement Network (PATS-UIENet). This network contains a Direct Signal Transmission Estimation Steam (D-Stream), a Backscatter Signal Transmission Estimation Steam (B-Stream) and an Ambient Light Estimation Stream (A-Stream), which are used to explicitly estimate the parameters of the revised IFM [2]. To address the challenge of the lack of labeled real-world underwater images, we further adopt an IFM-inspired semi-supervised learning framework, which consists of a bi-directional supervised scheme and an unsupervised scheme. Compared to the supervised or unsupervised method, the PATS-UIENet can be better trained using this framework with both the labeled and unlabeled real-world images while the generalization of the model trained is thus stronger, due to the complementary action of the two schemes.
III-A Revised IFM
As pointed out by Akkaynak and Treibitz [2], the original IFM [1] ignored the dependencies of the backscatter coefficient on the ambient light and the optical property of water bodies. They assumed that the attenuation coefficients of the direct signal and the backscattered signal are the same. In this case, the original IFM [1] cannot describe the degradation process of underwater images adequately.
Akkaynak and Treibitz [2] further proposed a revised IFM, which can be expressed as:
$I^{c}(x)=J^{c}(x)e^{-\beta_{c}^{D}(\mathbf{D})\cdot z}+\left(1-e^{-\beta_{c}^{B}(\mathbf{B})\cdot z}\right)A^{c},$ (2)
where $\textbf{D}=\{z,\rho,E,S_{c},\beta\}$ and $\textbf{B}=\{E,S_{c},b,\beta\}$ are two sets of parameters which affect the attenuation factors of the direct signal $\beta_{c}^{D}$ and the backscatter signal $\beta_{c}^{B}$, respectively, $z$ represents the distance between the scene and the camera, $\rho$ is the reflectance spectrum of the object, $E$ is the ambient light at a certain distance, $S_{c}$ is the camera response function, $b$ is the beam scattering coefficient and $\beta$ is the beam attenuation coefficient. The above formula can be written in the more compact form as:
$I^{c}(x)=J^{c}(x)t_{D}^{c}(x,\textbf{D})+\left(1-t_{B}^{c}(x,\textbf{B})\right)A^{c}.$ (3)
It has been demonstrated that the attenuation factors of the direct signal and the backscatter signal are different while the attenuation factor of the backscatter signal is affected by the ambient light more severely [2]. For more details, please refer to [2].
III-B PATS-UIENet
As illustrated in Fig. 2, the PATS-UIENet contains two encoder-decoder style streams with the same structure, i.e., D-stream and B-stream, and a Transformer-based A-stream. The three streams are used to estimate the direct signal transmission map, the backscatter signal transmission map and the ambient light, respectively. Since the red channel usually encounters more severe attenuation than the blue and green channels [1, 48], the utilization of this channel is key to color restoration. Therefore, a simple Red Channel Tuner (RCT) module is designed to adaptively emphasize the red channel before an image is fed into the transmission estimation steams.
The output of the first encoder block $Enc\_1$ in the B-stream is processed by Patchify and the result is sent to the A-stream. A Residual Communication Module (RCM) connects the two blocks of the encoder of the B-stream and the A-stream at the same level, to fulfill the bi-directional feature exchange. The two sets of transmission maps estimated using the D-stream and B-stream, respectively, and the ambient light estimated using the A-stream are fed into the revised IFM [2]. As a result, the enhanced underwater image is produced.
Figure 2: The architecture of the proposed PATS-UIENet, which comprises three individual streams, namely, D-Stream, B-Stream and A-Stream, to estimate the degradation parameters of the revised IFM [2].
Red Chanel Tuner (RCT). Inspired by the channel attention mechanism [49], we propose a compact RCT module in order to adaptively emphasize the information contained in the red channel. This module first uses a convolutional layer to extract basic features from the input image, and then use the Global Average Pooling (GAP) to obtain a global representation. A fully-connected layer and the Sigmoid activation function are further used to transform this representation to a tuning weight. The weight is finally used to scale the red channel of the image.
D-Stream. According to Eq. (3), the transmission map varies at different pixel locations. In this case, the network should be able to generate fine-grained representations. Hence, we build the D-Stream using a CNN-based encoder-decoder network. The encoder comprises five consecutive blocks, denoted as $Enc\_1$ to $Enc\_5$, and the Max-pooling operation is applied to each block. As a result, multi-scale feature maps are derived using the encoder. Given $Enc\_5$ is used as the bottleneck, the decoder contains four blocks, denoted as $Dec\_1$ to $Dec\_4$, symmetrically. Bilinear interpolation is used to upsample the feature maps within each block. Both the encoder and decoder blocks share the same structure, which consists of two $Conv$-$BN$-$ReLU$ units. Skip connection is used to pass the feature maps produced by an encoder block to the corresponding decoder block at the same level, which is useful for restoring the fine-grained spatial structure. The output of the last decoder block is fed into a convolutional layer. A set of transmission maps are produced in terms of different color channels.
B-Stream. Similarly, the B-stream is also designed as a CNN-based encoder-decoder network with the same architecture as the D-stream, motivated by Eq. (3). Unlike the D-stream, however, the output of the first encoder block $Enc\_1$ in the B-stream undergoes the Patchify processing, which includes both the flattening and linear projection operations, and the result is then passed to the A-stream. Besides, a Residual Communication Module (RCM) connects the $Enc\_2$ to $Enc\_5$ in the B-stream with the $Trans\_2$ to $Trans\_5$ in the A-stream at the same level, respectively.
A-Stream. The ambient light describes the luminance of the scene. Since this parameter does not encode the information of a specific object, it can be treated as homogeneous. We are motivated to design the A-stream using a set of Transformer [50] blocks for the sake of leveraging the self-attention mechanism to extract the global information. Five blocks, denoted as $Trans\_1$ to $Trans\_5$, are comprised of this stream. Each block contains a self-attention layer and a two-layer Multi-layer Perceptron (MLP) unit, in which two residual connections are used separately.
Since the input of the Transformer is a sequence of visual tokens, the feature maps of $Enc\_1$ are processed using Patchify. Considering that the estimation of the ambient light does not require local characteristics and the quadratic complexity of self-attention is unavoidable, we downsample these maps by a factor of 16 to reduce the computational demand. The result is a set of tokens. A special token, namely, the Ambient token, is also adopted, to specifically learn the ambient light parameter. All the tokens are fed into $Trans\_1$. The Ambient token produced by the last block $Trans\_5$ is sent to a linear layer. Finally, the estimation of the ambient light is achieved.
Residual Communication Module (RCM). According to the revised IFM [2], the degradation can be considered as the outcome caused by both the transmission and the ambient light. In this situation, the backscatter signal transmission estimation steam and the ambient light estimation stream will learn some common features. Therefore, the feature exchange between both the streams is likely to boost the training of each stream. Recently, the complementary action of CNNs and Transformers has been shown [51]. We are inspired to propose the RCM for the sake of bridging the two streams. Specifically, the tokens produced by a block in the A-stream are folded into a set of feature maps while the feature maps generated by a block of the encoder of the B-stream are downsampled to the shape of these maps. Then both sets of maps are concatenated along the channel dimension and are fed into a convolutional layer. Finally, the resultant maps are split into two sets along the channel dimension. Each set is added with the original feature maps of the related stream.
III-C IFM-Inspired Semi-supervised Learning Framework
To overcome the challenge of insufficient real-world underwater images, we adopt a semi-supervised learning framework (see Fig. 3), inspired by the revised IFM [2] theory. This framework includes a bi-directional supervised scheme and an unsupervised scheme. The bi-directional scheme contains a forward-enhancement module, which learns from reference images and uses the estimated parameters of the revised IFM to perform UIE, and a backward-degradation module, which uses the estimated parameters to degrade the reference images to their real-world degraded counterparts. As a result, the bi-directional supervised scheme can learn a more accurate estimation of the parameters using limited real-world data than that learned using a supervised method with a large number of synthetic data. Also, the generalization of the model trained is stronger than that trained using a supervised method with limited real-world data. On the other hand, the unsupervised scheme exploits the unlabeled real-world underwater images that we collect. A second set of images with different degrees of degradation are generated from these images based on the revised IFM [2]. Two sets of parameters can be estimated using both sets of images, respectively. Each set is utilized as the reference for the other set because their IFMs are related.
Figure 3: The proposed semi-supervised learning framework, which comprises a bi-directional supervised learning scheme and an unsupervised learning scheme, used for training our PATS-UIENet.
Bi-directional Supervised Learning. Given a set of labeled training data, we first use the estimated parameters to obtain enhanced results $\hat{J}^{c}(x)$. Then we perform the forward-enhancement learning module by minimizing the following loss function:
$\mathcal{L}_{fwd}=||(J^{c}(x)-\hat{J}^{c}(x))||_{2}^{2}.$ (4)
Due to the ill-posed nature of the revised IFM, however, only $\mathcal{L}_{fwd}$ may be insufficient for ensuring the validity of the estimated parameters. Therefore, we propose a backward-degradation learning module which degrades clear reference images using the estimated parameters to their degraded counterparts $\hat{I}^{c}(x)$. The supervision over the backward-degradation is achieved by minimizing the loss function:
$\mathcal{L}_{bwd}=||(I^{c}(x)-\hat{I}^{c}(x))||_{2}^{2}.$ (5)
Inspired by the Retinex [13] theory, we further apply Gaussian blur to input images in order to derive the hint of the global ambient light, which is able to boost the training of the A-stream. Given the ambient light $\hat{A}^{c}$ estimated using the A-stream, we minimize the following loss function:
$\mathcal{L}_{A\text{-}sup}=||(I^{c}(x)*G-\hat{A}^{c}||_{2}^{2},$ (6)
where $G$ stands for a Gaussian kernel and * is the convolution operation. Finally, the bi-directional supervised learning scheme is conducted as:
$\mathcal{L}_{sup}=\mathcal{L}_{fwd}+\lambda_{1}\mathcal{L}_{bwd}+\lambda_{2}\mathcal{L}_{A\text{-}sup},$ (7)
where $\lambda_{1}$ and $\lambda_{2}$ are used to balance different loss functions. Compared with the supervised methods [28, 22, 23], this scheme exploits the limited labeled underwater images better.
Unsupervised Learning. To leverage unlabeled underwater images, generative networks [30, 19] were used. However, they were difficult to train and produced unstable results. Since unsupervised learning methods [32, 33] are able to directly learn from degraded images without requiring reference images, we are motivated to adopt an unsupervised learning scheme on top of the revised IFM [2].
Given an unlabeled real-world degraded image $I^{c}_{1}(x)$, the revised IFM can be expressed as:
$I^{c}_{1}(x)=J^{c}_{1}(x)t_{D1}^{c}(x)+(1-t_{B1}^{c}(x))A^{c}_{1}.$ (8)
A new degraded image $I^{c}_{2}(x)$ can be purposely derived according to:
$I^{c}_{2}(x)=\alpha I^{c}_{1}(x)+(1-\alpha)(1-t_{B1}^{c}(x))A^{c}_{1},$ (9)
where $\alpha\in(0,1)$ is a controlled factor which decides the degradation extent. Substituting $I^{c}_{1}(x)$ in Eq. (8) into Eq. (9), $I^{c}_{2}(x)$ can be then expressed as:
$I^{c}_{2}(x)=J^{c}_{1}(x)(\alpha t_{D1}^{c}(x))+(1-t_{B1}^{c}(x))A^{c}_{1}.$ (10)
Eq. (10) also satisfies the revised IFM but presents a more severe degradation than the $I^{c}_{1}(x)$, generated from the same underlying clear image $J^{c}_{1}(x)$, because $\alpha t_{D1}^{c}(x)<t_{D1}^{c}(x)$ always holds true. Therefore, the parameters $\hat{t}_{D2}^{c}(x)$ and $\hat{t}_{B2}^{c}(x)$ estimated using the PATS-UIENet from image $I^{c}_{2}(x)$ will be close to $\alpha\hat{t}_{D1}^{c}(x)$ and $\hat{t}_{B1}^{c}(x)$, respectively, where $\hat{t}_{D1}^{c}(x)$ and $\hat{t}_{B1}^{c}(x)$ are the parameters estimated from $I^{c}_{1}(x)$. Correspondingly, two loss functions are defined as:
$\mathcal{L}_{T}=||\hat{t}_{D2}^{c}(x)-\alpha\hat{t}_{D1}^{c}(x)||_{2}^{2},$ (11)
$\mathcal{L}_{A\text{-}unsup}=||\hat{t}_{B1}^{c}(x)-\hat{t}_{B2}^{c}(x)||_{2}^{2}.$ (12)
Due to the lack of reference images, we use a non-reference loss function [33] based on the Gray-world prior [52] to constrain the enhancement of image $\hat{J}^{c}_{1}(x)$, which is expressed as:
$\mathcal{L}_{gw}=\displaystyle\sum_{c\in\{R,G,B\}}||E(J^{c}_{1}(x))-0.5||_{2}^{2},$ (13)
where $E(J^{c}_{1}(x))$ denotes the average of a specific channel $c$. Finally, we define the loss function of the unsupervised learning scheme as:
$\mathcal{L}_{unsup}=\mathcal{L}_{T}+\mathcal{L}_{A\text{-}unsup}+\lambda_{3}\mathcal{L}_{gw},$ (14)
where $\lambda_{3}$ is a weighting factor.
Semi-supervised Learning. On top of both the bi-directional supervised learning and unsupervised learning schemes, the proposed semi-supervised learning framework can be formulated as:
$\mathcal{L}_{semi\text{-}sup}=\mathcal{L}_{sup}+\lambda_{unsup}\mathcal{L}_{unsup},$ (15)
where $\lambda_{unsup}$ is used to balance the two schemes.
IV Experimental Settings
In this section, we will briefly introduce the baselines, data sets, evaluation metrics and implementation details utilized in our experiments.
IV-A Baselines
We compared the proposed method with four prior-based approaches which were developed on top of the IFM [1], including DCP [3], UDCP [4], GDCP [5] and DHL [10], four prior-based methods which did not consider the IFM, including Histogram Equalization (HE) [9], Retinex [13], MMLE [14] and HLRP [15], and eight deep learning approaches, including WaterNet [22], UColor [23], FUnIE-GAN [20], USUIR [24], PUIE-Net (MP) [25], PUIE-Net (MC) [25], U-Transformer [21] and URanker [26].
IV-B Data Sets
We conducted a series of experiments using five data sets, including four publicly available real-world underwater data sets, i.e., UIEB [22], SUIM-E [27], RUIE [53] and SQUID [10], and a synthetic data set, namely, EUVP [20]. The UIEB data set originally contains 890 pairs of degraded images, in which each reference image was selected from the results produced by applying 12 algorithms to a degraded image. Ten pairs of UIEB [22] images were discarded because the degraded images in these pairs had been included in the rest pairs. In total, we only used 880 pairs of UIEB images.
The SUIM-E [27] data set comprises 1,525 pairs of degraded underwater images and a testing set of 110 labeled images. Three subsets are comprised of the RUIE [53] data set, including UIQS, UCCS and UHTS, which include 3,630, 300 and 300 unlabeled images, respectively. The SQUID [10] data set comprises 114 underwater images captured using a stereo camera, divided into the Michmoret, Katzaa, Nachsholim and Satil subsets. We utilized 2,185 images from the Underwater Scenes subset of the EUVP [20] data set.
Two different PATS-UIENet models were trained for the real-world and synthetic testing images, referred to as $PATS\text{-}UIENet_{Real}$ and $PATS\text{-}UIENet_{Syn}$, respectively. Regarding both the models, we randomly selected 720 pairs from the 880 pairs of UIEB [22] images and 2,000 pairs from the Underwater Scenes subset of the EUVP [20] data set, respectively, which were used as the training images in the supervised learning scheme. We collected 1,934 unlabeled real-world underwater images, which cover diverse underwater scenes, and used these images in the unsupervised learning scheme for both the models.
Among the rest of the UIEB [22] data set, 80 pairs of images were randomly selected as the validation set, while the remaining pairs were used as the testing set which is referred to as Test-U80. The UIEB data set also consists of 60 challenging images without references, which were used as the second testing set, referred to as Test-C60. A third testing set was obtained from the testing set of the SUIM-E data set [27], which contains 110 pairs of labeled images and is named as Test-S110. For the RUIE [53] data set, the UCCS was utilized as the fourth testing set, denoted as Test-UCCS. Fifty-three SQUID [10] images captured using the camera at the right hand side were randomly selected and were used as the fifth testing set, namely, Test-R53. The remaining 185 images in the Underwater Scenes subset of the EUVP [20] data set were used as the sixth testing set, referred to as Test-Scenes.
IV-C Evaluation Metrics
For the labeled testing data, we used both Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [54] to assess the quality of an enhanced image with regard to the reference image. Regarding the unlabeled testing data, we used the Underwater Image Quality Measure (UIQM) [55] and Underwater Color Image Quality Evaluation (UCIQE) [56] metrics. When the SQUID [10] data set was tested, the Average Reproduction Angular Error ($\bar{\psi}$) [10] was used to evaluate the quality of color restoration. Since transmission maps are related to the scene distance, we used the Pearson Correlation Coefficient (PCC) calculated between the estimated transmission maps and the ground-truth depth map to assess the performance of transmission estimation [10].
IV-D Implementation Details
We implemented the proposed method using Pytorch and conducted experiments on Ubuntu 20.04 with a GeForce RTX 3090 graphics processing unit. During the training process, the images were resized to the resolution of $256\times 256$ pixels. The number of filters in the RCT was set to 16. For the $Enc\_1$ to $Dec\_1$ in the D-stream and B-stream, the numbers of filters were set to 64, 128, 256, 512, 512, 256, 128, 64 and 64 in turn. Regarding the A-stream, the dimension of each token was 384 and the number of attention heads in the MHSA was 6. We trained the PATS-UIENet using the AdamW [57] optimizer. The learning rate and batch size were set to 1e-4 and 12, respectively. Four weighting factors, including $\lambda_{1},\lambda_{2},\lambda_{3}$ and $\lambda_{unsup}$, were set to 0.001, 0.005, 10 and 0.001, respectively. We first trained the PATS-UIENet using the bi-directional supervised scheme for 3,000 iterations as a warm-up stage. Then the semi-supervised scheme was used to train it for 150,000 iterations. When the semi-supervised learning process was carried out, in particular, the unsupervised scheme was performed for 30 iterations after the bi-directional supervised scheme was conducted for every 120 iterations.
V Experimental Results
In this section, we will report the results obtained in the underwater image enhancement, transmission estimation and color restoration experiments and the ablation study.
V-A Underwater Image Enhancement
We evaluated the proposed PATS-UIENet along with 16 baselines using two full-reference quantitative metrics and two non-reference quantitative metrics. A qualitative analysis was also performed. In this subsection, the results obtained in the three experiments will be reported.
V-A1 Full-Reference Quantitative Evaluation
Since the Test-U80, Test-S110, and Test-Scenes testing sets contain reference images, a full-reference quantitative evaluation was conducted on the proposed method and the 16 baselines using the PSNR and SSIM metrics across these testing sets. The results are shown in Table I. As can be seen, our PATS-UIENet achieved the best result in terms of both the metrics no matter which testing set was used. Compared with the prior-based baselines using IFM [1], such as DCP [3], UDCP [4], GDCP [5] and DHL [8], our method showed a significant improvement.
TABLE I: The full-reference quantitative evaluation of the proposed PATS-UIENet and sixteen baselines on three real-world testing sets. The best and second best results are highlighted in the Red Bold and Blue Italic fonts, respectively.
Method Test-U80 Test-S110 Test-Scenes
[…]
DCP 12.48 63.34 14.52 67.38 18.09 68.95
UDCP 9.37 33.62 10.04 34.28 14.76 56.43
GDCP 12.46 67.41 14.67 71.29 14.11 63.50
DHL 14.49 62.81 14.96 69.68 15.16 53.92
HE 16.48 77.77 15.37 74.76 13.61 62.82
Retinex 10.38 61.25 8.29 53.44 9.50 55.56
MMLE 18.56 76.21 17.32 77.39 14.89 62.23
HLRP 13.56 22.76 12.55 29.24 12.12 18.21
WaterNet 17.02 70.27 18.80 75.19 25.84 84.02
UColor 20.93 85.19 20.23 84.23 24.31 81.40
FUnIE-GAN 15.31 59.68 16.89 66.32 25.11 83.16
USUIR 16.92 68.69 17.54 76.58 15.14 66.51
PUIE-Net (MP) 22.26 88.89 21.85 89.19 21.85 73.47
PUIE-Net (MC) 22.05 88.80 21.72 89.20 20.85 73.10
U-Transformer 20.98 72.76 19.92 68.25 24.09 79.98
URanker 21.87 85.75 21.27 87.43 22.46 82.35
PATS-UIENet (Ours) 23.01 89.89 22.48 90.26 27.20 88.30
V-A2 Non-reference Quantitative Evaluation
We used two non-reference metrics, i.e., UIQM [55] and UCIQE [56], to evaluate the performance of our method and sixteen baselines on the Test-U80, Test-S110, Test-Scenes, Test-C60, Test-UCCS and Test-R53 testing sets. The results are shown in Tables II. It can be seen that the performance of our method was lower than the best counterpart but was comparable to most baselines. As demonstrated in [22, 27, 35, 26], the UIQM and UCIQE metrics are not always consistent with the quality of images processed by UIE methods. These findings may explain the relatively lower UIQM and UCIQE values produced by our method.
TABLE II: The non-reference quantitative evaluation of the proposed PATS-UIENet and sixteen baselines on six real-world testing sets. The best and second best results are highlighted in the Red Bold and Blue Italic fonts, respectively.
Method Test-U80 Test-S110 Test-Scenes Test-C60 Test-UCCS Test-R53
[…]
DCP 1.35 0.51 1.36 0.57 1.61 0.57 0.56 0.51 0.03 0.44 -2.05 0.42
UDCP 2.69 0.59 1.23 0.60 2.80 0.60 1.24 0.54 1.83 0.57 0.76 0.59
GDCP 2.17 0.56 2.12 0.64 1.89 0.65 1.49 0.60 1.24 0.56 5.50 0.50
DHL 3.19 0.64 2.35 0.66 3.54 0.66 1.85 0.64 2.98 0.63 0.22 0.64
HE 3.58 0.67 2.97 0.68 4.27 0.67 2.32 0.65 3.93 0.67 1.71 0.64
Retinex 2.33 0.48 0.85 0.45 3.36 0.47 1.21 0.46 2.03 0.47 -2.03 0.51
MMLE 3.89 0.58 2.72 0.60 3.60 0.61 2.77 0.57 4.06 0.57 3.18 0.55
HLRP 3.71 0.63 2.88 0.65 3.85 0.66 1.98 0.65 3.89 0.64 1.98 0.65
WaterNet 2.11 0.50 1.47 0.54 3.67 0.57 1.28 0.51 1.25 0.45 0.09 0.45
UColor 3.44 0.60 2.30 0.63 3.64 0.57 2.08 0.58 3.22 0.55 1.28 0.54
FUnIE-GAN 3.73 0.50 3.59 0.57 4.43 0.58 4.50 0.51 3.47 0.49 0.19 0.42
USUIR 4.10 0.63 2.84 0.65 3.88 0.65 3.12 0.62 4.19 0.60 3.31 0.60
PUIE-Net (MP) 3.42 0.56 2.25 0.59 3.54 0.57 1.82 0.55 4.03 0.52 1.56 0.51
PUIE-Net (MC) 3.31 0.56 2.22 0.59 3.11 0.57 1.78 0.55 3.82 0.52 1.30 0.49
U-Transformer 4.37 0.55 3.68 0.56 4.54 0.55 4.19 0.53 4.19 0.53 2.43 0.53
URanker 3.64 0.57 2.55 0.59 3.68 0.57 2.36 0.56 4.17 0.56 2.25 0.56
PATS-UIENet (Ours) 3.76 0.57 2.47 0.60 3.70 0.57 2.24 0.56 3.58 0.55 2.55 0.55
V-A3 Qualitative Analysis
The visual results generated by the sixteen baselines and our method on six testing sets are shown in Figs. 4 - 9, respectively. It can be seen that some methods with the higher UIQM [55] or UCIQE [56] score did not produce visually pleased results. For example, MMLE [14] increased the contrast but produced images with pale colors (see Fig. 4). FUnIE-GAN [20] could not completely remove the color cast in degraded images (see Fig. 7). HE[9], HLRP[15], and USUIR[24] introduced artifacts that reduced the visual quality (see Fig. 8). In contrast, our method not only effectively removed the color cast from degraded images but also significantly improved the visibility of the images. It is demonstrated that the images that our method enhanced still exhibit the satisfactory visual quality even if the UIQM or UCIQE value that our method generated is lower than that produced by the baselines. This observation is consistent with the findings presented in the literature [22, 27, 35, 26].
Figure 4: The results produced by 16 baselines and our method in terms of a degraded image in the Test-U80 testing set.
Figure 5: The results produced by 16 baselines and our method in terms of a degraded image in the Test-S110 testing set.
Figure 6: The results produced by 16 baselines and our method in terms of a degraded image in the Test-Scenes testing set.
Figure 7: The results produced by 16 baselines and our method in terms of a degraded image in the Test-C60 testing set.
Figure 8: The results produced by 16 baselines and our method in terms of a degraded image in the Test-UCCS testing set.
Figure 9: The results produced by 16 baselines and our method in terms of a degraded image in the Test-R53 testing set.
V-B Transmission Estimation and Color Restoration
The quantitative evaluation of our methods and four prior-based baselines for transmission estimation and color restoration is reported in Table III. It can be seen that our method achieved the superior, or at least the comparable, performance to that of the four prior-based baselines on four different subsets of Test-R53. Specifically, our method outperformed the baselines with a large margin on both the Michmoret and Katzaa subsets, no matter which metric was considered. It should be noted that none of these methods produced a PCC value higher than 0.12 on the Satil subset. The inferior results may be attributed to the special scenes contained in this subset [35]. As shown in Fig. 10, our method was able to perform transmission estimation and color restoration by directly learning from the limited real-world underwater images.
Figure 10: Each group shows a degraded image and the red channels of five transmission maps (top) and five enhanced images (bottom) obtained using four prior-based baselines and our method.
TABLE III: The quantitative evaluation of four prior-based baselines and our method on the four subsets of Test-R53, where PCC and $\bar{\psi}$ are used for transmission estimation and color restoration, respectively.
Method Michmoret Katzaa Nachsholim Satil
[…]
DCP -0.16 34.63 0.00 35.47 -0.26 34.75 0.03 36.21
UDCP -0.54 36.97 -0.15 40.43 0.06 38.80 0.12 51.45
GDCP 0.29 33.76 0.24 34.47 -0.04 34.28 -0.03 35.31
DHL -0.07 32.10 0.10 35.50 0.36 35.92 0.11 34.82
Ours 0.61 11.68 0.32 9.47 -0.38 16.87 0.02 10.36
V-C Ablation Study
To investigate the effect of different components of the PATS-UIENet, we conducted a series of ablation experiments. For simplicity, we only utilized Test-U80 and the Michmorest subset in Test-R53.
V-C1 Effect of the Semi-supervised Learning Framework
To further validate the effect of semi-supervised learning framework, we provide a quantitative comparison among the supervised learning, bi-directional supervised learning, and semi-supervised learning frameworks. As reported in Table IV, either the supervised learning framework or the bi-directional supervised learning framework normally generated the lower metric values than those produced by the proposed semi-supervised learning framework, in terms of the PSNR, SSIM, PCC and $\bar{\psi}$ metrics. It has been suggested that our network could be better trained using the semi-supervised learning framework with both the labeled and unlabeled real-world images, which improved the enhancement performance.
TABLE IV: The ablation study of different learning frameworks using Test-U80 and the Michmorest subset in Test-R53.
Learning Framework Test-U80 Michmoret
[…]
Sup 21.09 86.72 0.51 17.86
Bi-sup 22.06 88.95 0.65 15.59
Semi-sup (Ours) 23.01 89.89 0.61 11.68
V-C2 Effect of the RCT and the RCM
To examine the effect of different modules on the performance of our PATS-UIENet, we removed the RCT, the RCM or both of them. In this case, we derived three variants of the PATS-UIENet: one with both RCT and RCM removed, one with only RCT removed and one with only RCM removed, denoted as “Simplified”, “w/o RCT” and “w/o RCM”, respectively. The results produced by these variants and the PATS-UIENet are reported in Table V. It can be observed that the PATS-UIENet with the entire components normally achieved the best performance in terms of the PSNR, SSIM, PCC and $\bar{\psi}$ metrics. Although the utilization of the RCT had a negative effect on the PSNR value (observed by comparing “Simplified” with “w/o RCM”), the performance of transmission estimation was increased. It was indicated that the adaptability of the red channel increased the robustness and generalization of the model trained.
TABLE V: The ablation study of different modules on Test-U80 and the Michmorest subset in Test-R53.
Variant Test-U80 Michmoret
[…]
Simplified 22.18 88.56 0.59 13.35
w/o RCT 22.71 89.43 0.60 11.42
w/o RCM 21.84 88.98 0.61 14.05
PATS-UIENet 23.01 89.89 0.61 11.68
V-C3 Effect of Different Network Depth
To evaluate the impact of the network depth on the performance of our PATS-UIENet, we changed the number of layers of the network from 4 to 3 and 5, individually. In total, we derived two variants of the PATS-UIENet with regard to different network depths. In terms of both the variants, the numbers of filters at each block were set to 64, 128, 256, 256, 128, 64 and 64 and 64, 128, 256, 512, 1024, 1024, 512, 256, 128, 64 and 64, respectively. As shown in Table VI, the proposed PATS-UIENet which comprised four layers produced the more promising results compared with the two variants.
TABLE VI: The ablation study of different network depths on Test-U80 and the Michmorest subset in Test-R53.
Network Depth Test-U80 Michmoret
[…]
3 Layers 22.43 89.40 0.61 13.29
5 Layers 22.57 89.45 0.62 9.41
4 Layers (Ours) 23.01 89.89 0.61 11.68
V-C4 Effect of the Revised IFM
To further investigate the effect of the revised IFM [2] on UIE, we used the outputs of the B-Stream and A-Stream of the proposed PATS-UIENet as the transmission map and the ambient light data, respectively, required by the original IFM [1], as shown in Eq. (1). In this case, the results obtained using the original IFM are compared with those produced by the revised IFM in Table VII. As can be seen, the revised IFM achieved the superior performance to the original IFM in terms of the PSNR, SSIM and $\bar{\psi}$ metrics. It is indicated that the revised IFM models the underwater degradation process better than the orignal IFM.
TABLE VII: The comparison between the original IFM [1] and the revised IFM [2] on Test-U80 and the Michmorest subset in Test-R53.
Physical Model Test-U80 Michmoret
[…]
Original IFM [1] 21.93 88.24 0.69 16.24
Revised IFM [2] 23.01 89.89 0.61 11.68
VI Conclusion
In this paper, we introduced a novel Physics-Aware Triple-Stream Underwater Image Enhancement Network, namely, PATS-UIENet, which explicitly estimates the degradation parameters of the revised IFM, for the UIE task. To overcome the challenge of insufficient data, we also adopted an IFM-inspired semi-supervised learning framework, comprising a bi-directional supervised scheme and an unsupervised scheme. Due to the complementary action of the two schemes, the PATS-UIENet can be better trained using this framework with both the labeled and unlabeled real-world images while the generalization of the model is stronger, compared to supervised and unsupervised methods. To our knowledge, this study made the first effort to jointly exploit the physics-aware deep network and the IFM-inspired semi-supervised learning technique for the UIE task. The proposed method performed better than, or at least comparably to, sixteen baselines on six underwater testing sets in different evaluations. The promising results should be due to the fact that our method is able to not only estimate degradation parameters but also learn the characteristics of diverse underwater scenes.